OpenStudy (my4babiesandme):

write a quadretic equation in the variable x having the given numbers as solutions type the equation in standard form, ax^2+bx+c=0 -sqrt6,8sqrt6

OpenStudy (anonymous):

\[(x + \sqrt{6})(x - 8\sqrt{6}) = 0\]

OpenStudy (anonymous):

\[x^2 -8x\sqrt{6} + x\sqrt{6} - 48 = 0\] \[\text{The answer is:}\] \[x^2 - 7x\sqrt{6} - 48 = 0\]

OpenStudy (my4babiesandme):

thank you